# MADNESS Deblender

### v1

@@ 4.3.1 True residuals
For the baseline, we use the true color value obtained from the catalog ($C^{T}$). Figure 12a shows the variation of the true residual ($\delta C^{T}$) mathematically given by
@@
[…]
The distribution of $\delta C^{T}$ shows a subtle skew as the contamination increases, and in different bins of true color, we observe a slight bias being introduced in the bins at both ends of the plot. This pattern of bias is expected because of the statistics of true galaxy color - i.e., a galaxy with a very low value for g-r color (for example, $-0.5$) is most likely to be blended with a galaxy of higher color value which will cause the observed value of g-r color to increase and vice versa.
@@ 4.3.2 Blending residuals
[…]
@@ 4.4 Errors in morphology and shapes of the deblended galaxies
Figure 13: Distribution of scores for morphology metrics: SSIM and cosine similarity ($S_{\text{cos}}$) for the deblended models of MADNESS, scarlet, and VAE-deblender in the r-band of galaxies from the TestFields. For both metrics, values closer to 1 imply better performance.
(a) Distribution of e1 and $e_{2}$ residuals for the entire TestFields dataset.
(b) Distribution of e1 and $e_{2}$ residuals for galaxies satisfying $\kappa_{r}>0.01$.
Figure 14: Comparison of the ellipticities from second order moments among MADNESS, scarlet, and VAE-deblender. The ellipticity residuals are shown for (a) entire TestFields dataset and (b) slightly blended galaxies with $\kappa_{r}>0.01$, which represents at least $1\%$ flux contribution from neighboring sources.
@@
In this section, we evaluate the morphology of the reconstructed galaxies. Besides the SSIM metric, introduced in Section 3.2, we also use the cosine similarity $S_{\text{cos}}$, both of which give us an idea of the similarity in structural information in the images. Mathematically, $S_{\text{cos}}$ is given by
@@
Both SSIM and $S_{\text{cos}}$ metrics range from $-1$ (anti-correlation) to $+1$ (exact prediction). To evaluate the deblenders, we compute the two metrics between the simulated galaxy and the deblended models separately in each band. In Figure 13 we compare SSIM and cosine similarity of the reconstructions in the r-band. We observe that the performance of the VAE-deblender is very close to that of scarlet and MADNESS performs better than both scarlet and the VAE-deblender. To be more quantitative, with galaxies from TestFields satisfying $\kappa_{r}\leq 5$, the average value for $S_{\text{cos}}$, for example, is $0.994$, $0.991$, and $0.987$ for MADNESS, scarlet, and VAE-deblender, respectively.
To evaluate the shapes of the reconstructions, we use a definition of ellipticity $e$ that is commonly used in weak lensing analyses [^11] See Part 3, Fig. 2 in Schneider (2006), for example.:
@@
We compare the ellipticity residuals defined as the difference between the ellipticities $e_{1}$ or $e_{2}$ of the deblended models and the values obtained from isolated noiseless galaxy images. The distribution of ellipticity residuals in Figure 14a shows that for the entire TestFields dataset MADNESS performs better than both scarlet and the VAE-deblender. For example, the standard deviation of the $e_{1}$ residuals is $0.069$, $0.084$, and $0.104$ for MADNESS, scarlet, and VAE-deblender respectively. Figure 14b shows the distribution of residuals upon selecting only blended galaxies with contamination $\kappa_{\textit{r}}\geq 0.01$. The standard deviation for $e_{1}$ residuals with the contamination cut is $0.087$ for MADNESS and $0.103$ for scarlet. The relatively better performance of MADNESS compared to scarlet is in agreement with our results from Section 4.2 where we argued that slightly blended galaxies are likely to be isolated for scarlet after PSF deconvolution hence allowing more robust reconstruction for scarlet.
@@ 5 Towards real data
Before we apply MADNESS to real data, a few important steps require consideration. MADNESS must be combined with a detection algorithm as it does not include a detection step. In our work, we restricted ourselves to the ground truth positions to evaluate the efficiency of the deblender but in reality, a significant fraction of these blends are likely to go unrecognized and the galaxies may not be perfectly centered. Moreover, multi-band processing will also add subtleties like coaddition and World Coordinate Systems (WCS) which may vary across bands. Further work is required to systematically evaluate the biases introduced during the deblending process when MADNESS is combined with a detection algorithm. In fact, the final value of log-prior in equation 15 might provide us with valuable insight to flag potentially unrecognized blends, and space-based surveys can be useful to obtain a prior detection which adds to the motivation for the joint-processing of data.
In this work, we have only considered bulge plus disk models and for each galaxy, all parameters except the magnitudes and PSF are assumed to be constant across all bands, which is not necessarily the case for real data (Casura et al., 2022; Häußler et al., 2022). This warrants further studies with more realistic simulations consisting of complex galaxy morphologies. Additionally, the algorithm needs to be flexible enough to properly handle stars. This can be done either by including stars in the training dataset of the VAE itself so that the generative model can simulate both stars and galaxies or by having a hierarchical framework that includes a dedicated star galaxy separation algorithm (see Hansen et al., 2022).
Finally, our work can also be extended into a hierarchical Bayesian framework to incorporate PSF information extracted by other science pipelines. Following the same ideas as Lanusse et al. (2021), MADNESS can be trained to predict the PSF-deconvolved galaxies and reconvolved with externally obtained PSF before computing the likelihood. Furthermore, the lower level of blending in space-based telescopes due to smaller PSFs implies that joint processing data can significantly help to deblend ground-based observations. Arcelin et al. (2020) already showed that VAEs can be used for multi-instrument processing by concatenating data as different channels to the CNNs of the VAE. In particular, the authors show that the results of deblending were significantly improved when combining data from Euclid with Rubin. Since MADNESS also makes use of a VAE, we expect this to generalize and hope to obtain a significant boost by multi-instrument processing.
@@ 6 Conclusions
Blending is expected to be a major source of systematic errors in the next generation of cosmological surveys. The complexity of the problem given the wide range of morphologies and colors of astrophysical objects makes it an ideal domain for the application of ML techniques that have already proven to be extremely efficient for image generation and processing. In this context, we developed a new deblending algorithm called MADNESS, which combines two generative models, namely the VAE and NF, in a way so that we can obtain the MAP solution for the inverse problem of deblending.
@@
We trained and tested our algorithm with galaxies expressed as a sum of bulge and disk models taken from the CatSim catalog. The galaxies were simulated for a 10-year LSST survey using GalSim with the help of BTK (Mendoza et al., in prep.), which allows us to easily control the number and position of galaxies. We first trained a $\beta-$VAE as a generative model for isolated galaxies and then trained an NF to learn the underlying aggregate posterior distribution in the VAE latent space. The NF gives us access to the log-prior term in the VAE latent space and the decoder provides the mapping from the latent space to the image space, thus enabling us to compute the log-likelihood term needed to obtain the MAP solution. Combining the two models we perform a gradient descent in the VAE latent space to obtain a MAP estimate. Typically, deep learning-based deblenders require the training data to be representative of the real data. While this is also true for MADNESS, the optimization step makes our algorithm less susceptible to this than traditional feed-forward networks and at the same time, it makes our algorithm more interpretable.
@@
To enable faster convergence of our algorithm, we retrained the encoder to obtain an initial position to start the optimization. We call this new encoder the VAE-deblender because the encoder is trained to find the latent space representation of the central galaxy in a blended postage stamp, thereby deblending it from neighboring sources, similar to the network proposed by Arcelin et al. (2020). Although the VAE-deblender is very efficient in itself, the results thus obtained are subject to the black box nature of the algorithm and lack explicit flux optimization. In our algorithm, we take the latent space representations obtained from the VAE-deblender as the initial point and perform a gradient descent in the VAE latent space to obtain the MAP solution.
@@
We tested our algorithm for flux reconstruction with aperture-photometry, the g-r color of the deblended models, and morphology using SSIM and the pixel cosine similarity as metrics. As a sanity check, we compared our results to the initial deblending solutions of the VAE-deblender, and observed a significant improvement for all metrics in the final MAP solution. We also compared MADNESS to the state-of-the-art deblender scarlet. Looking at the results in the r-bandpass filter, we observed that the VAE-deblender was already able to closely match the performance of scarlet. However, when we look at the results from other filters in Appendix D and E of LSST we notice a significant drop in performance of both MADNESS and the VAE-deblender compared to scarlet in the u-band, especially in terms of morphology. This is because scarlet predicts a joint morphology for all bands and the high-fidelity bandpass filters (such as r, i, etc.) allowing better reconstructions.
@@
In the case of MADNESS and the VAE-deblender, although the VAE learns correlations between different bands, the predicted morphology can still vary between the different channels of VAE output. When obtaining the MAP estimate, MADNESS does try to minimize the data likelihood so we expect the overall flux to be consistent even in low-SNR bands. However, reconstructing the morphology is more difficult as the edges of the galaxies will have even lower SNR than the central brighter regions.
@@
Our algorithm can easily be expanded to a hierarchical framework to incorporate physical information such as PSF along with noise. Furthermore, it can also be extended into a multi-instrument approach by concatenating data as new channels to the VAE input. Arcelin et al. (2020) already showed that doing so can improve the results of deblending and we expect those results to generalize. Such algorithms that are capable of combining data at the pixel level are likely to play a key role in the upcoming golden era of observational cosmology where we will have data from several Stage IV ground and space-based surveys.
@@ Acknowledgements.
This paper has undergone internal review in the LSST Dark Energy Science Collaboration. The authors would like to kindly thank the internal reviewers: Patricia Burchat and Cyrille Doux. Their comments, feedback, and suggestions helped us greatly improve the paper. We also would like to thank François Lanusse, Ismael Mendoza, Peter Melchior, Fred Moolekamp, James J. Buchanan and Tianqing Zhang for comments and discussions throughout this work. B.B. led the project, wrote the code, trained the neural networks, performed the analysis, and wrote the initial draft of the paper. E.A., A.B., A.G., J.L., and C.R. supervised the project, reviewed the code, and provided feedback on the draft. This project has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 945304 – COFUND AI4theSciences hosted by PSL University, and Agence Nationale de la Recherche grant ANR-19-CE23-0024 — AstroDeep. We gratefully acknowledge support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing and data-processing resources needed for this work. The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy, the National Science Foundation, and the LSST Corporation in the United States. DESC uses resources of the IN2P3 Computing Center (CC-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515. The authors acknowledge the use of the following libraries: numpy, astropy, TensorFlow, TensorFlow-Probability, scarlet, Blending-Toolkit, GalSim, SurveyCodex, SourceExtractor, and matplotlib.
Appendix A SNR computation
Following the LSST documentation [^12] https://smtn-002.lsst.io/, SNR is defined as
$\text{SNR}=\frac{C}{\sqrt{C+(\frac{B}{g}+\sigma^{2}_{\text{instr}})n_{\text{eff}}}},$ (24)
where $C$ is the total source counts, $B$ is sky background counts per pixel, $\sigma_{\text{instr}}$ is the instrumental noise per pixel (all in ADU), $g$ is the gain, and $n_{\text{eff}}$ refers to the source footprint. We discard the $\sigma_{\text{instr}}$ term as our simulations do not take it into account. We use SurveyCodex to obtain the $B$ and $C$ parameters for each galaxy and set the value of gain to $1$ for consistency with our BTK simulations. Finally, we evaluate $n_{\text{eff}}$ as
@@ Appendix A SNR computation
where $\text{FWHM}_{\text{eff}}$ is the effective full width at half maximum and $P$ refers to the pixel scale, which for LSST is $0.2$″per pixel. We compute the $\text{FWHM}_{\text{eff}}$ as a convolution of PSF FWHM and galaxy FWHM ($\text{FHWM}_{\text{gal}}$) given by
@@
where $a_{\text{hlr}}$, and $b_{\text{hlr}}$ are the semi-major and semi-minor axis of either the bulge or the disk, whichever is greater.
Appendix B scarlet Configuration

### v2

@@ 4.3.1 True residuals
For the baseline, we used the true color value obtained from the catalog ($C^{T}$). Figure 10a shows the variation of the true residual ($\delta C^{T}$) mathematically given by
@@
where $m\in\{\text{no deblending, {MADNESS}\@, {scarlet}\@}\}$ and correspondingly $C_{m}$ refers to the g-r color measured on TestFields in the three cases: without deblending, deblending with MADNESS, and deblending with scarlet. We see in Figure 10a that even for low contamination, errors are induced by the noise in images, but they are symmetrical around zero so there is no bias. As the contamination increases, the symmetry breaks and the uncertainty increases. However, both deblenders efficiently remove the bias and reduce the errors introduced by blending.
The distribution of $\delta C^{T}$ shows a subtle skew as the contamination increases, and in different bins of true color, we observe a slight bias being introduced in the bins at both ends of the plot. This pattern of bias is expected because of the statistics of true galaxy color - i.e., a galaxy with a very low value for g-r color (for example, 0.5-0.5) is most likely to be blended with a galaxy of higher color value, which will cause the observed value of g-r color to increase and vice versa.
@@ 4.3.2 Blending residuals
where once again $m\in\{\text{no deblending, {MADNESS}\@, {scarlet}\@}\}$. The distributions of $\delta C^{B}$ in Figure 10b back our conclusions from Section 4.3.1 that blending leads to a bias in our measurement of the color. Moreover, with $\delta C^{B}$ we can directly compare the deblenders. We see in Figure 10b that although both MADNESS and scarlet reduce the error and bias in color, MADNESS reconstructs the g-r color more accurately.
@@ 4.4 Errors in morphology and shapes of the deblended galaxies
Figure 11: Distribution of scores for morphology metrics: SSIM and cosine similarity ($S_{\text{cos}}$) for the deblended models of MADNESS, scarlet, and VAE-deblender in the r-band of galaxies from the TestFields. For both metrics, values closer to 1 imply a better performance.
Figure 12: Comparison of the ellipticities from second-order moments among MADNESS, scarlet, and VAE-deblender. The ellipticity residuals are shown for (a) entire TestFields dataset and (b) slightly blended galaxies with $\kappa_{r}>0.01$, which represents an at least $1\%$ flux contribution from neighboring sources.
@@
In this section, we evaluate the morphology of the reconstructed galaxies. Besides the SSIM metric, introduced in Section 3.2, we also use the cosine similarity $S_{\text{cos}}$, both of which give an idea of the similarity in structural information in the images. Mathematically, $S_{\text{cos}}$ is given by
@@
Both SSIM and $S_{\text{cos}}$ metrics range from 1-1 (anti-correlation) to +1+1 (exact prediction). To evaluate the deblenders, we computed the two metrics between the simulated galaxy and the deblended models separately in each band. In Figure 11, we compare the SSIM and cosine similarity of the reconstructions in the r-band. We observe that the performance of the VAE-deblender is very close to that of scarlet, and MADNESS performs better than both scarlet and the VAE-deblender. To be more quantitative, with galaxies from TestFields satisfying $\kappa_{r}\leq 5$, the average value for $S_{\text{cos}}$, for example, is $0.994$, $0.991$, and $0.987$ for MADNESS, scarlet, and VAE-deblender, respectively.
To evaluate the shapes of the reconstructions, we used a definition of ellipticity $e$ that is commonly used in weak lensing analyses [^10] See Part 3, Fig. 2 in Schneider (2006), for example.:
@@
We compare the ellipticity residuals defined as the difference between the ellipticities $e_{1}$ or $e_{2}$ of the deblended models and the values obtained from isolated noiseless galaxy images. The distribution of ellipticity residuals in Figure 12a shows that for the entire TestFields dataset MADNESS performs better than both scarlet and the VAE-deblender. For example, the standard deviation of the $e_{1}$ residuals is $0.069$, $0.084$, and $0.104$ for MADNESS, scarlet, and VAE-deblender, respectively. Figure 12b shows the distribution of residuals upon selecting only blended galaxies with contamination $\kappa_{\textit{r}}\geq 0.01$. The standard deviation for $e_{1}$ residuals with the contamination cut is $0.087$ for MADNESS and $0.103$ for scarlet. The relatively better performance of MADNESS compared to scarlet is in agreement with our results from Section 4.2.
@@ 5 Towards real data
Before we apply MADNESS to real data, a few important steps require consideration. MADNESS must be combined with a detection algorithm, as it does not include a detection step. In our work, we restricted ourselves to the ground-truth positions to evaluate the efficiency of the deblender, but in reality, a significant fraction of these blends are likely to go unrecognized, and the galaxies may not be perfectly centered. Additionally, it might be necessary to retain the networks with lower S/N objects to handle false detections. Moreover, multi-band processing will also add subtleties such as coaddition and World Coordinate Systems (WCS), which may vary across bands. Further work is required to systematically evaluate the biases introduced during the deblending process when MADNESS is combined with a detection algorithm. In fact, the final value of log-prior in Equation 10 may provide valuable insight to flag potentially unrecognized blends, and space-based surveys can be useful for obtaining prior detections, which adds to the motivation for the joint-processing of data.
In this work, we have only considered bulge plus disk models and for each galaxy. Except for the magnitudes and PSF, all galaxy parameters related to shapes and sizes are assumed to be constant across all bands. This assumption is not necessarily the case for real data (Casura et al., 2022; Häußler et al., 2022), and this warrants further follow-up studies with more realistic simulations consisting of complex galaxy morphologies. Additionally, the algorithm needs to be flexible enough to handle stars properly. This can be done either by including stars in the training dataset of the VAE itself so that the generative model can simulate both stars and galaxies or by having a hierarchical framework that includes a dedicated star galaxy separation algorithm (see Hansen et al., 2022).
Finally, our work can also be extended into a hierarchical Bayesian framework to incorporate PSF information extracted by other science pipelines. Following the same ideas as Lanusse et al. (2021), MADNESS can be trained to predict the PSF-deconvolved galaxies and reconvolved with externally obtained PSF before computing the likelihood. Furthermore, the lower level of blending in space-based telescopes due to smaller PSFs implies that joint processing data can significantly help to deblend ground-based observations. Arcelin et al. (2020) already showed that VAEs can be used for multi-instrument processing by concatenating data as different channels to the input layer of the VAE. In particular, the authors show that the results of deblending were significantly improved when combining data from Euclid with Rubin. Since MADNESS also makes use of a VAE, we expect this to generalize and hope to obtain a significant boost by multi-instrument processing.
@@ 6 Conclusions
Blending is expected to be a major source of systematic errors in the next generation of cosmological surveys. The complexity of the problem, given the wide range of morphologies and colors of astrophysical objects, makes it an ideal domain for the application of ML techniques that have already proven to be extremely efficient for image generation and processing. In this context, we developed a new deblending algorithm called MADNESS, which combines two generative models, namely the VAE and normalizing flow, in a way so that we can obtain the MAP solution for the inverse problem of deblending.
@@
We trained and tested our algorithm with galaxies expressed as a sum of bulge and disk models taken from the CatSim catalog. The galaxies were simulated for a 10-year LSST survey using GalSim with the help of BTK (Mendoza et al., 2025), which allows one to easily control the number and position of galaxies. We first trained a $\beta$-VAE as a generative model for isolated galaxies and then trained a normalizing flow to learn the underlying aggregate posterior distribution in the VAE latent space. The normalizing flow gives access to the log-prior term in the VAE latent space, and the decoder provides the mapping from the latent space to the image space, thus enabling the computation of the log-likelihood term needed to obtain the MAP solution. Combining the two models, we perform a gradient descent in the VAE latent space to obtain a MAP estimate. Typically, deep learning-based deblenders require training data to be representative of real data. While this is also true for MADNESS, the optimization step makes our algorithm less susceptible to this than traditional feed-forward networks and at the same time, it makes our algorithm more interpretable.
@@
To enable faster convergence of our algorithm, we retrained the encoder to obtain an initial position to start the optimization. We call this new encoder the VAE-deblender because the encoder is trained to find the latent space representation of the central galaxy in a blended postage stamp, thereby deblending it from neighboring sources, similar to the network proposed by Arcelin et al. (2020). Although the VAE-deblender is very efficient in itself, the results thus obtained are subject to the black-box nature of the algorithm and lack explicit flux optimization. In our algorithm, we take the latent space representations obtained from the VAE-deblender as the initial point and perform a gradient descent in the VAE latent space to obtain the MAP solution.
@@
We tested our algorithm for flux reconstruction with aperture-photometry, the g-r color of the deblended models, and morphology using SSIM and the pixel cosine similarity as metrics. As a sanity check, we compared our results to the initial deblending solutions of the VAE-deblender, and observed a significant improvement for all metrics in the final MAP solution. We also compared MADNESS to the state-of-the-art deblender scarlet. Looking at the results in the r-bandpass filter, we observed that the VAE-deblender was already able to closely match the performance of scarlet. However, when we look at the results from other filters in Appendix F and G of LSST, we notice a significant drop in performance of both MADNESS and the VAE-deblender compared to scarlet in the u-band, especially in terms of morphology. This is because scarlet predicts a joint morphology for all bands and the high-fidelity bandpass filters (such as r, i, etc.) allow better reconstructions.
@@
In the case of MADNESS and the VAE-deblender, although the VAE learns correlations between different bands, the predicted morphology can still vary between the different channels of VAE output. When obtaining the MAP estimate, MADNESS tries to minimize the data likelihood, so we expect the overall flux to be consistent even in low-S/N bands. However, reconstructing the morphology is more difficult as the edges of the galaxies will have even lower S/N than the central brighter regions.
@@
Our algorithm can be easily expanded to a hierarchical framework to incorporate physical information, such as PSF, along with noise. Furthermore, it can also be extended into a multi-instrument approach by concatenating data as new channels to the VAE input. Arcelin et al. (2020) already showed that doing so can improve the deblending results, and we expect them to generalize. Such algorithms that are capable of combining data at the pixel level are likely to play a key role in the upcoming golden era of observational cosmology, where we will have data from several stage IV ground- and space-based surveys.
@@ Acknowledgements.
This paper has undergone internal review in the LSST Dark Energy Science Collaboration. The authors would like to kindly thank the internal reviewers, Patricia Burchat and Cyrille Doux. Their comments, feedback, and suggestions helped us greatly improve the paper. We also would like to thank François Lanusse, Ismael Mendoza, Peter Melchior, Fred Moolekamp, James J. Buchanan and Tianqing Zhang for comments and discussions throughout this work. B.B. led the project, wrote the code, trained the neural networks, performed the analysis, and wrote the initial draft of the paper. E.A., A.B., A.G., J.L., and C.R. supervised the project, reviewed the code, and provided feedback on the draft. This project has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 945304 – COFUND AI4theSciences hosted by PSL University, and Agence Nationale de la Recherche grant ANR-19-CE23-0024 — AstroDeep. We gratefully acknowledge support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing and data-processing resources needed for this work. The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy and the LSST Discovery Alliance in the United States. DESC uses resources of the IN2P3 Computing Center (CC-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515. The authors acknowledge the use of the following libraries: numpy, astropy, TensorFlow, TensorFlow-Probability, scarlet, Blending-Toolkit, GalSim, SurveyCodex, SourceExtractor, and matplotlib.
Appendix A S/N computation
Following the LSST documentation [^11] https://smtn-002.lsst.io/, S/N is defined as
$\text{S/N}=\frac{C}{\sqrt{C+(\frac{B}{g}+\sigma^{2}_{\text{instr}})n_{\text{eff}}}},$ (19)
where $C$ is the total source counts, $B$ is sky background counts per pixel, $\sigma_{\text{instr}}$ is the instrumental noise per pixel (all in ADU), $g$ is the gain, and $n_{\text{eff}}$ refers to the source footprint. We discard the $\sigma_{\text{instr}}$ term as our simulations do not take it into account. We used SurveyCodex to obtain the $B$ and $C$ parameters for each galaxy and set the value of gain to $1$ for consistency with our BTK simulations. Finally, we evaluate $n_{\text{eff}}$ as
@@ Appendix A S/N computation
where $\text{FWHM}_{\text{eff}}$ is the effective full width at half maximum and $P$ refers to the pixel scale, which for LSST is 0.20.2″per pixel. We compute the $\text{FWHM}_{\text{eff}}$ as a convolution of PSF FWHM and galaxy FWHM ($\text{FHWM}_{\text{gal}}$) given by
@@
where $a_{\text{hlr}}$ and $b_{\text{hlr}}$ are the semi-major and semi-minor axes of either the bulge or the disk, whichever is greater.
Appendix B Architectures and training of networks
B.1 VAE
Figure 14: Variational autoencoder architecture. The conv1, conv2, conv3, and conv4 layers refer to the four convolutional layers of the encoder, all of which use a kernel size of 5 and a stride of 2. Following the convolutional layers are the dense layers d1 and d2 that predict the parameters of the latent space distribution. The probabilistic layer p1 uses the output of layer d2 to construct the latent space $z$, which is a multivariate normal distribution of 16 dimensions. Finally, the decoder takes as input a sample from the latent space distribution and applies a set of dense layers d3 and d4, followed by transposed convolutional layers (convT1, convT2, convT3, and convT4) also with kernel size 5 and stride 2; Finally, the output of convT4 is cropped to match the size of the input.
In our implementation, we used the TensorFlow  [^12] https://www.tensorflow.org/ framework and TensorFlow-Probability  [^13] https://www.tensorflow.org/probability for the probabilistic layers. As shown in Figure 14, the encoder has 4 convolutional layers with 64, 128, 256, and 128 filters. Each layer uses a convolution kernel of size 5 and a stride of 2. These layers are followed by 2 dense layers that predict the latent space distribution. Each layer in the encoder is followed by PReLU activation functions except the final dense layer, which is left without an activation function as it predicts the parameters for a multivariate normal distribution layer that describes the $16$-dimensional latent space distribution. The latent space is followed by the decoder, which also has two dense layers and 4 transposed convolutional layers with 128, 96, 64, and 6 filters. The first 3 layers use a kernel of size 5 and stride 2, while the final layer has a kernel of size 3 and stride 1. We used the PReLU activation function in all the layers except the final layer, where we applied a ReLU activation function to prevent negative predictions of flux. Finally, the output is cropped to match the size of the input. The encoder and decoder have around 6.76.7M and $3.5\rm{M}$ trainable parameters, respectively. The architecture of the encoder is chosen to be more complex because it is later trained as a deblender to obtain the initial point for gradient descent, as described in Section 3.4.
During training, we optimized the loss function in Equation 2, where $\beta$, the KL divergence coefficient, is set to $0.01$. By using $\beta<1$, we can improve the reconstruction fidelity, but this also implies that the learned aggregate posterior differs from the prior. An alternative approach could be to remove the KL term entirely. However, the lack of regularization will lead to a more complex latent space distribution, as there is no incentive for the VAE to follow the ELBO-prior. The added complexity of the latent space would make it more difficult to perform a gradient descent in the VAE latent space for the MAP estimate. After testing the $\beta$-VAE for several values of the hyperparameter $\beta$, we chose the value of $0.01$ as no significant improvement was observed by lowering the value further.
The coefficient $\alpha$ is initialized at $1$ and is decreased by $0.02$ every epoch until it reaches a value of 0 after $50$ epochs. The SSIM is computed on normalized images where each band of input and output is divided by the maximum pixel value in that band, so all values are in the range $[0,1]$. Since the $\operatorname{\mathbb{S}}$ is equally weighted among all the bands, the network is forced to learn the features in all the filters from the very beginning.
We used the Adam optimizer (Kingma & Ba 2017) with an initial learning rate of $10^{-5}$, along with a learning rate scheduler that decreases the learning rate by a factor of $2.5$ every $40$ epochs. Additionally, since we performed only a linear normalization of our data and some galaxies are extremely bright compared to others, we introduce a clip value of $0.1$ on the gradients obtained during the optimization to stabilize the training of our network. After the initial warm-up phase is complete, we dropped the SSIM term from the loss function and restarted the training with a learning rate of $10^{-5}$ for a maximum of $200$ epochs with an early stopping patience of $20$ so that the network stops training if the loss has not decreased over $20$ iterations. Even during this training phase, we keep the same learning rate scheduler and clip value.
B.2 Normalizing flows
We used TensorFlow-Probability to create bijective autoregressive networks with 2 hidden layers containing 32 units each followed by tanh activation functions, to be used within MAF layers. After an initial batchnorm layer, we use 6 sets of MAF followed by permutation and batchnorm layers, resulting in a model with a total of 15 93615\,936 trainable parameters (see Figure 15). We used the Adam optimizer and started with a learning rate of $10^{-4}$, clip value of $0.01$, and the same setting for the learning rate scheduler and patience as the VAE, for a maximum of $200$ epochs.
Figure 15: Schematic of the training of the normalizing flow. The encoder is the pre-trained network illustrated in Figure 14 and the normalizing flow consists of 6 bijective transformations $f_{\varphi_{i}}$ where the $i^{th}$ transformation is parameterized by $\varphi_{i}$, which represents the weights of the MAF layer containing (32, 32) units with tanh activations followed by a permutation layer. Additionally, a batchnorm layer is added as the first layer within $f_{\varphi_{1}}$.
B.3 VAE-deblender
We ran this training for a maximum of $200$ epochs with the Adam optimizer and a learning rate of $10^{-5}$. Other parameters such as the clip value, learning rate scheduler, and patience were kept the same as the original VAE.
Appendix C Comparison between u- and r-band
Since the training of the VAE uses the likelihood obtained from pixel-wise predictions, the quality of reconstruction depends upon the S/N of the galaxies. In Figure C, we show a comparison between the quality of simulations of our trained model in the u and r bands for the same set of galaxies. This comparison demonstrates that the model is able to better learn the features in the higher S/N band compared to the lower S/N one. So we expect poor deblending reconstructions in the u-band as shown in Appendix F and G.
Figure 16: Comparison of simulations for the same galaxies between the u-band shown in panel (a) and r-band shown in panel (b). The color bar in each image shows the normalized electron counts in pixels.
Appendix D scarlet configuration